TPC QA pre run 2025 cosmics

 
Summary of pre run cosmics  5/8/2025

The information is in chronological order

-----------

Mail from 3/29

 As mentioned at today's 10 o'clock meeting the noise has again shown up with the same features as earlier in the week

The sequence was the following
good up until run 87017  8:26
then some misc activity in the hall

87013 was first pedestal run after the access.
87026 first cosmic run \ reasonable good but signs of noise apparent in RDObyte plot
by run 87039 really bad (2014):
stayed bad until morning when I requested a pedestal run at the meeting.
At the first run after the pedestal run it was ok. and still is now (16:30 Saturday).


I look a bit more and it seems it is not just noise in the adc spectrum, but gives rise to to increased #clusters as observed in some of the #clusters per sector (inner).This is unlikely a problem for lasers and cosmics.

We need to think about this and discuss how to isolate the problem.
Observations:
What is common for sector 3 and 4 as the seems to go bad together.
--- ideas RDO power supplies sharing
---  gating grid noise.
     Is it not true that for iTPC there is no timebin dependent pedestal so strange noise patterns could disturb values uploaded ?

Is there evidence from the data from the pedestal runs if it is a shift in pedestal, and increase width (i.e noise level).
It is spread out across all timebins, and we see it across all pads and rows.
A particular feature seems to be a peak in the middle of the time spectrum.






Summary from FV  on 5/2/2025

I summarizes runs for the cosmics we took and documented when problem occurred.
excel file

When I look at when things went bad- two cases but not noticed for long time it seems to have happended after a long break ours magnet trip magnet taken down. The bad noise only appeared after such period.
I one of the cases there was a pedestal run done, and initial runs had only little nosie but subseqy=uent more
(, but 3rd run after that would always be still bad.
fter a (on 29 after a ped run I request it was good until we stopped.

It seems to me that this is a warmup-up issue. Is it poiisble that more power is delivered to 3+4?
or something like that just a thought.








response from Tonko after investigation 5/8/25
---------

here seems to be a shift in pedestals of 3-4-5 ADC counts for
2 different but specific FEEs: 1 in Sector 3, 1 in Sector 4.
All timebins of 1 SAMPA chip (this 1/2 a FEE) jump up and
stay up. No idea why. Could be malfunctioning SAMPA.
The regulators are shared between the 2 SAMPAs so that's
not the reason. In any case, not much we can do.
 
You can see the odd shape of the cluster charge step with a sharp
narrow peak in the middle: the cluster finder treats this as 1 biiiig
cluster spanning 400ish timebins and calculates the mean timebin at 
~half of the range i.e. 200.
 
E.g. run 26082041.
 
I'd like to let it be like this for a day of the next run (I assume a few days
of cosmics) and I'll take another look. But the way out is to simply 
kill those FEEs. The reason I don't want to do it just now because
I am curious and I want to write some code which compares pedestal
runs "in vivo" looking for this pedestal jump.

FV/ you can also see in the tpcclustertimewidth that there are many assigned 30 i.e largest values
and matches what Tonko say.